# Mitigating the Human-Robot Domain Discrepancy in Visual Pre-training for Robotic Manipulation

### v1

@@ Abstract
Learning generalizable visual dynamic representation across different embodied environments is crucial for real-world robotic manipulation. As the scale and diversity of robot demonstration data are limited, recent works have turned to large-scale pre-training using human data. However, the morphological differences between humans and robots introduce a significant human-robot domain discrepancy, challenging the generalization of these human-data pre-trained models to downstream manipulation tasks. To address this, we propose a novel adaptation paradigm that utilizes readily available paired human-robot video data to bridge the discrepancy. Following this paradigm, our method exploits a human-robot contrastive alignment loss to align the semantics of human and robot videos, adapting pre-trained models to the robotic domain in a parameter-efficient manner. The experiments demonstrate significant improvements on 25 tasks across three different benchmarks, where the single-task, language-conditioned multi-task settings are covered, and two different pre-trained models are evaluated. On the large RLBench benchmark, our adaptation method achieves an average improvement of $8.9\%$ in success rate over the pre-trained R3M model [1] across multiple tasks. We will release the code and models upon acceptance. project: https://jiaming-zhou.github.io/projects/HumanRobotAlign
@@ 1 Introduction
The objective of robotic manipulation [2; 3; 4; 5; 6; 7; 8; 1] is to equip robots with diverse skills applicable in real-life scenarios. This task is notably challenging as it necessitates modeling the sequential semantics of visual observations to accurately predict the subsequent end-effector actions. Given the limited availability of robot demonstration data, several efforts [1; 9; 10; 11; 12] have aimed at pre-training visual models on large-scale human data [13; 14; 15; 16; 17], e.g., Ego4D [14]. These pre-trained models will serve as frozen visual backbones to learn manipulation policies for downstream tasks. The core intuition [18] driving this is the dynamic similarity between human-object interactions and robot action sequences when performing similar tasks. This similarity facilitates the application of visual models, learned from human data, to downstream robotic manipulation tasks.
@@
Nonetheless, the significant variance in visual elements (e.g., scenes, views, objects), as well as the intrinsic morphological differences between humans and robots, results in a vast domain discrepancy [19; 20] between human and robot data. Therefore, as depicted at the top of Figure 1, the existing paradigm cannot ensure a successful transfer of models pre-trained on extensive human data to downstream manipulation tasks. Recently, two different approaches have been proposed to solve this problem. One [10; 21] introduces manipulation-oriented proxy tasks when pre-training models on human data. For example, Huo et al. [10] fine-tuned multiple perceptual tasks (e.g., hand detection) on top of pre-trained encoders to adapt the learned representation for learning downstream manipulation policies. However, due to the lack of robot data during fine-tuning, this approach cannot explicitly alleviate the domain discrepancy, and the manipulation-oriented tasks are hard to define on human data. The second approach [22; 23] integrates learnable modules into pre-trained models and fine-tunes the models on downstream robot data to learn more discriminative representation. However, this approach requires customizing the pre-trained models for each distinct environment, which reduces the versatility of the pre-trained models.
@@
In this work, we propose a new adaptation paradigm to mitigate the human-robot domain discrepancy while maintaining the versatility of pre-trained models for robotic manipulation, as depicted at the bottom of Figure 1. To achieve this, we use a small amount of paired human-robot demonstration data, which serves as a bridge. This data has two key characteristics. First, each pair of human and robot demonstrations has well-aligned semantics. Second, although the used robot demonstrations are visually different from downstream robot data, the domain discrepancy between them is reduced due to their homogeneous robot structure (see Figure 1). The core idea is to adapt the pre-trained models using this data so that the semantics in robot demonstrations captured by the adapted models are consistent with those in the paired human demonstrations. This paradigm effectively alleviates the discrepancy between pre-training and downstream robot data while retaining the versatility of the adapted models, eliminating the need for customization for each downstream environment. Following our new paradigm, we propose an efficient Human-Robot Semantic Alignment (HR-Align) method, which adapts the pre-trained models by incorporating the parameter-efficient adapter module [24; 25; 22] into the models. To guide the adaptation towards reducing the domain discrepancy, our HR-Align method employs a human-robot contrastive alignment loss to improve the semantic consistency between features of the paired human and robot videos.
@@
  • We highlight the human-robot domain discrepancy in visual pre-training for robotic manipulation, and provide a new adaptation paradigm that simultaneously alleviates the domain discrepancy and maintains the versatility of pre-trained models;
@@
  • We evaluate the effectiveness of our method in three different environments, covering 7 single-tasks and 18 language-conditioned multi-tasks, as well as pre-trained models with different pre-training methodologies.
Figure 1: Design principle of our proposed adaptation paradigm for visual pre-training. Top: The previous paradigm applied frozen models pre-trained on human data to downstream robot data, suffering from large domain gaps. Bottom: Our adaptation paradigm leverages the paired human-robot demonstrations as bridges to adapt the pre-trained models, thus mitigating the semantics gap between human and robot data. (Best viewed in color.)
@@ 2 Related Works
Visual-based robotic manipulation. The rapid evolution in the field of robotic manipulation [2; 3; 4; 5; 6; 7; 8; 1] in recent years is noteworthy. Initial investigations [4; 26; 27; 28] developed visuo-motor policies focused on reinforcement learning [29; 30; 31]. This allowed robots to learn from their environments, although high sample complexity and difficult reward engineering were involved. More recent studies [6; 7; 8; 12; 32; 33] have geared towards learning policies by imitating actions in expert demonstrations, which has yielded promising results. In addition, the community of robotic manipulation has been further propelled by the introduction of new benchmarks. Transitioning from benchmarks with limited scales and diversity [34; 35; 36] to larger scale benchmarks [28; 37; 2; 38] with increased object, scene, and task variations (e.g., RLBench [28] and CALVIN [37]). For example, on the RLBench benchmark [28], many research works [6; 7; 12; 32; 39; 33] have been proposed to solve the robotic manipulation tasks.
@@
Robotic visual pre-training. To efficiently teach robots manipulation skills for real-life scenarios, it is crucial to pre-train models on large-scale data. This ensures that the generalizable visual dynamic representation across different environments is effectively learned. Drawing inspiration from the representation learning [40; 41; 42; 43; 44; 45] in computer vision, there have been several endeavors [46; 47; 2; 48; 49; 50] to accomplish this goal in robotic manipulation. The RT-series [46; 47; 2], for instance, developed vision-language-action models using large-scale data that encompassed out-of-domain data from the internet and self-gathered robotic demonstrations. With a 3D-instruction tuning dataset stemming from multiple sources of robot data, 3D-VLA [48] trained a 3D-based generative world model capable of generating actions for robot manipulation. RoboFlamingo [49] learned to generate robot actions by fine-tuning pre-existing vision-language foundation models on robot demonstrations. However, the collection of large-scale robot demonstration data proved costly, and unresolved issues remain regarding the different formats between data from different embodiments [51; 50; 2].
@@
In parallel, inspired by human action understanding [52; 53; 54; 55; 56; 57; 58; 59; 60], several studies [1; 9; 11; 12; 61; 62; 63; 64] have proposed methods to learn a generalizable representation from large-scale human datasets, such as Ego4D [14] and Kinetics [65]. These methods are based on the assumption that the dynamics of human-object interactions in these datasets are analogous to those of robots performing tasks. For instance, the R3M [1] model was pre-trained on the Ego4D dataset using time-contrastive learning and video-language alignment techniques. Similarly, the MVP [9] model employed masked modeling to learn visual dynamics, and the data4robotics [11] performed a dataset-centric analysis of robotic pre-training, contributing valuable insights to the field. Despite these advances, significant challenges remain due to the inherent morphological differences between humans and robots. These differences lead to a human-robot domain discrepancy [10; 23; 22; 66; 18; 67], which limits the generalizability of models trained on human data when applied to downstream robotic tasks.
@@
Human-robot domain transfer. Some works [10; 22; 23] in human-data pre-training have noticed the human-robot domain gap problem and have developed two main strategies to address it. The first is to design more effective proxy tasks to learn semantics that can be better transferred to downstream robotic manipulation when pre-training models on the human data. For example, Huo et al. [10] proposed human-oriented multi-task fine-tuning on top of pre-trained models to encode meaningful semantics for manipulation. However, the manipulation-oriented proxy tasks are difficult to define, and this implicit fine-tuning without accessing robot data is not effective in mitigating the human-robot domain discrepancy. The second strategy [22; 23] involves directly fine-tuning the pre-trained models on downstream manipulation data to bridge the semantic gap between human and robot data encountered during pre-training. For example, Sharma et al. [22] proposed to insert lightweight adapters into pre-trained models, aiming to reduce the performance gap between frozen pre-trained representation and full end-to-end fine-tuning on downstream data. However, this strategy requires custom modifications to the pre-trained models for each specific environment, which can compromise the models’ generalizability.
@@ 3 Human-Robot Semantic Alignment Method
Considering the large domain discrepancy between the pre-trained human data and downstream robotic manipulation data in the existing pre-training paradigm, this work proposes a new adaptation paradigm. This paradigm leverages the semantic-aligned human-robot demonstration data to guide the adaptation of existing pre-trained models, effectively enhancing their generalization capabilities across various downstream tasks. Driven by this idea, we propose an efficient Human-Robot Semantic Alignment method.
@@ 3.1 Problem formulation
Assuming we have a dataset $\mathcal{D}$ encompassing $N$ human-robot video pairs, i.e., $\mathcal{D}=\{H_{i},R_{i},L_{i}\}_{i=1}^{N}$, where $L_{i}$ is a sentence describing what task is being performed in the video pair, $H_{i}$ denotes the $i$-th human video demonstration, and $R_{i}$ denotes the $i$-th robot video demonstration that imitates the human actions in $H_{i}$ via teleoperation. Thus the dynamic semantics in each human-robot video pair are well-aligned. For learning the downstream manipulation policies, instead of using the robot’s raw image observations $I$ as state input, existing visual pre-training methods train visual encoder $\mathcal{F}$ on human data, and use the feature representation $\mathcal{F}(I)$ of robot’s observations as state input. In this work, to mitigate the human-robot domain discrepancy in the pre-trained representation of $\mathcal{F}$, we leverage the aforementioned paired human-robot video dataset $\mathcal{D}$ to adapt the encoder $\mathcal{F}$. In this way, the adapted model $\mathcal{T}$ can extract more effective state representations for learning policies in downstream robotic tasks.
@@ 3.2 Overview
Figure 2 provides an overall diagram of our Human-Robot Semantic Alignment (HR-Align) method. Taking the paired human and robot videos as input, the HR-Align method utilizes a frozen pre-trained model to obtain the feature representations of human video and robot video, respectively. However, since the used backbone model is pre-trained on human data, the semantics encoded by the extracted features of human and robot videos suffer from domain discrepancy. For the robot video, our HR-Align injects learnable adapter modules into the pre-trained model, so that the pre-trained model can be adapted to encode the dynamic semantics in the robot video. Besides, for each stream, the corresponding task description of the videos will serve as the query to extract task-aware semantics from the aforementioned feature representations. To guarantee the introduced adapter modules can be effectively trained for adapting the pre-trained model to the robot domain, the HR-Align leverages a contrastive alignment loss, which constraints the semantics in robot video encoded by the adapted model to align with the dynamic semantics in the paired human video.
3.3 Learning of HR-Align method
@@
where $\Upsilon$ denotes the frame sampling operation, and $h^{f},r^{f}\in\mathbb{R}^{\mathrm{\hat{T}\times\hat{H}\times\hat{W}\times\hat{C}}}$ are the frozen features extracted from the human video and robot video, respectively. As the backbone model $\mathcal{F}$ is well pre-trained on large-scale human data, the visual dynamic semantics in human video $H$ can be effectively captured. However, due to the significant human-robot domain gap, it is difficult to extract corresponding semantics from the paired robot video, using the model $\mathcal{F}$ that does not access any knowledge in the robot domain during pre-training.
@@
Adapted robot stream. To this end, our HR-Align method aims to efficiently adapt the pre-trained model on the robot videos, such that the newly adapted pre-trained model $\mathcal{T}$ can alleviate the domain discrepancy. To achieve this, inspired by existing parameter-efficient fine-tuning approaches [24; 25], we form the adapted pre-trained model $\mathcal{T}$ by inserting learnable adapter modules (Eq. (3)) into the pre-trained model $\mathcal{F}$. We formulate this as follows:
$r^{t}=\mathcal{T}(\Upsilon(R))=\mathcal{F}_{Adapter}(\Upsilon(R)),$ (2)
@@
where $r^{t}\in\mathbb{R}^{\mathrm{\hat{T}\times\hat{H}\times\hat{W}\times\hat{C}}}$ is the adapted robot video features. The adapter modules can be inserted into any position of the pre-trained model $\mathcal{F}$. Without loss of generality, given the output $r^{t,inter}\in\mathbb{R}^{\mathrm{\bar{T}\times\bar{H}\times\bar{W}\times\bar{C}}}$ of the intermediate layer of model $\mathcal{T}$, the adapter module can be applied as follows:
$r^{t,next}$ $=Adapter(r^{t,inter})$ (3)
$=r^{t,inter}+\text{Conv}_{up}(f(\text{Conv}_{down}(r^{t,inter}))),$
@@
where the residual output $r^{t,next}$ will be processed by subsequent layers of the model $\mathcal{T}$. Following existing practice [24; 25], our adapter module includes an activation function $f$, and two convolution layers (i.e., $\text{Conv}_{down}$ and $\text{Conv}_{up}$) used for feature down-projection and up-projection respectively.
@@
Task-aware feature modeling. To further enhance the feature representations of human and robot videos, i.e., the frozen human feature $h^{f}$, frozen robot feature $r^{f}$, and the adapted robot feature $r^{t}$, we use the feature of corresponding task description $L$ as the query to extract the task-aware semantics from these features. The feature of task description query $l\in\mathbb{R}^{\hat{C}}$ is obtained using a frozen BERT model [68; 69], with a learnable fully connected layer followed to match the channel dimension of the video features. The process is formulated as $l=Linear(Bert(L))$. Later, an attention-based aggregation is defined, where the task description feature $l$ serves as the query, and the video features in each stream serve as keys and values. Taking the reshaped video features $r^{t}\in\mathbb{R}^{\mathrm{(\hat{T}\cdot\hat{H}\cdot\hat{W})\times\hat{C}}}$ in the adapted robot stream as an example, the task-aware adapted robot feature $\bar{r}^{t}$ is defined as:
@@
For the frozen human and robot data stream, the task-aware frozen human feature $\bar{h}^{f}\in\mathbb{R}^{\hat{C}}$ and frozen robot feature $\bar{r}^{f}\in\mathbb{R}^{\hat{C}}$ can be obtained in the same way.
Human-Robot contrastive alignment. To mitigate the human-robot domain discrepancy in robotic pre-training, our key idea is to finetune the pre-trained model with a parameter-efficient design, where the semantic-aligned human-robot video pairs serve as bridges for the adaptation. For each human-robot video pair, the human-data pre-trained model can effectively encode the visual semantics in the human video. To guarantee the consistent semantics in the paired robot video can be captured by the adapted model, we propose a human-robot contrastive alignment loss to modulate the adaptation of the pre-trained model.
@@
In Eq. (6), $\mathcal{S}$ is a similarity function defined as $\mathcal{S}(x,y)=\exp(x^{T}y/\tau)$, where $\tau$ is a temperature factor. Under the constraint of the contrastive alignment loss, the learnable adapter modules enable the pre-trained model to capture robotic dynamics, thereby simultaneously mitigating the domain discrepancy and avoiding the need to customize models for each downstream environment.

### v3

@@ Abstract
Learning generalizable visual representations across different embodied environments is essential for effective robotic manipulation in real-world scenarios. However, the limited scale and diversity of robot demonstration data pose a significant challenge. Recent research has explored leveraging large-scale human activity data for pre-training, but the substantial morphological differences between humans and robots introduce a significant human-robot domain discrepancy, hindering the generalization of these models to downstream manipulation tasks. To overcome this, we propose a novel adaptation paradigm that leverages readily available paired human-robot video data to bridge the domain gap. Our method employs a human-robot contrastive alignment loss to align the semantics of human and robot videos, adapting pre-trained models to the robot domain in a parameter-efficient manner. Experiments on 20 simulated tasks across two different benchmarks and five real-world tasks demonstrate significant improvements. These results span both single-task and language-conditioned multi-task settings, evaluated using two different pre-trained models. Compared to existing pre-trained models, our adaptation method improves the average success rate by over $7\%$ across multiple tasks on both simulated benchmarks and real-world evaluations. Project: https://jiaming-zhou.github.io/projects/HumanRobotAlign
Figure 1: Left: The existing visual pre-training paradigm for robotic manipulation applied frozen models pre-trained on human data to downstream robot data, suffering from large domain gaps. Right: Our adaptation paradigm leverages the paired human-robot demonstrations as bridges to adapt the pre-trained models, thus mitigating the semantics gap between human and robot data. (Best viewed in color.)
@@ 1 Introduction
Robotic manipulation [63, 59, 35, 61, 10, 19, 11, 50] aims to enable robots to acquire a diverse range of skills applicable to real-world scenarios. However, this task is inherently challenging as it requires accurately modeling the sequential semantics of visual observations to predict the corresponding end-effector actions. Due to the scarcity of robot demonstration data, various approaches [50, 68, 29, 13, 36, 45, 24] have explored pre-training visual models on large-scale human activity datasets [57, 21, 20, 14, 12], such as Ego4D [21]. These pre-trained models are then used as frozen visual backbones to facilitate learning manipulation policies for downstream robotic tasks. The central intuition [1] underlying this paradigm is the dynamic similarity between human-object interactions and robot actions in similar tasks. This similarity facilitates the application of visual models, learned from human data, to downstream robotic manipulation tasks.
@@
Nonetheless, the intrinsic morphological differences between humans and robots lead to a substantial domain discrepancy between human and robot data [18, 3]. As depicted in the left part of Figure 1, the existing paradigm cannot ensure an effective transfer of models pre-trained on extensive human data to downstream manipulation tasks. Recently, two primary approaches have been proposed to solve this problem. The first [29, 2] focuses on introducing manipulation-oriented proxy tasks during pre-training on human data. For instance, Huo et al. [29] proposed training perceptual tasks (e.g., hand detection) on top of pre-trained models to adapt their representations for downstream robotic policies. However, this method lacks explicit access to robot data during training, making it difficult to directly mitigate the domain discrepancy. Moreover, defining manipulation-oriented tasks on human data remains a complex challenge. The second approach [58, 41] integrates learnable modules into pre-trained models and fine-tunes the models on downstream robot data to learn more discriminative representation. However, this approach requires customizing the pre-trained models for each distinct environment, which reduces the versatility of the pre-trained models. As a result, a dilemma arises of adapting human-data pre-trained representations to the robot domain while maintaining the model’s versatility.
@@
In this work, we propose a new adaptation paradigm, which mitigates the human-robot domain discrepancy while maintaining the versatility of pre-trained models for different downstream manipulation tasks, as depicted in the right part of Figure 1. To achieve this, we leverage a small amount of paired human-robot demonstration data as a bridge for adapting the pre-trained models. Such data, readily available in the community (e.g., the RH20T dataset [16]), exhibit two essential characteristics. First, each pair of human and robot demonstrations has well-aligned semantics. Second, although the used robot demonstrations are visually different from downstream robot data, the domain discrepancy between them is reduced due to their homogeneous robot structure. Leveraging the similar dynamic semantics within the paired human-robot demonstrations, the key insight motivating our adaptation is that the semantics in robot demonstrations, as captured by the adapted models, should be consistent with the pre-trained semantics in the paired human demonstrations. This paradigm effectively alleviates the discrepancy between pre-trains and downstream robot data while preserving the versatility of the adapted models, thereby eliminating the need for customization for each downstream environment. Building on this new paradigm, we introduce Human-Robot Semantic Alignment (HR-Align), a simple yet effective method for adapting pre-trained models. HR-Align incorporates a parameter-efficient adapter module [73, 51, 58] into the models to facilitate adaptation. To guide the adaptation towards reducing the domain discrepancy, our HR-Align method employs a human-robot contrastive alignment loss to ensure semantic consistency between the representations of paired human and robot data.
@@
  • We highlight the human-robot domain discrepancy problem in visual pre-training for robotic manipulation, and provide a new adaptation paradigm that simultaneously alleviates the domain discrepancy and maintains the versatility of pre-trained models;
@@
  • We demonstrate the effectiveness of our method on 20 simulated tasks across two benchmarks and five real-world tasks. Compared to existing pre-trained models, our method achieves over a $7\%$ improvement in average success rate across both single-task and multi-task settings.
@@ 2 Related Works
Visual-based robotic manipulation. Driven by the rapid evolution of computer vision [38, 25, 79, 78, 80, 28, 17], the developments of robotic manipulation [63, 59, 35, 61, 71, 66, 65, 67, 10, 19, 11, 50] in recent years are noteworthy. Initial investigations [35, 23, 42, 32] developed visuo-motor policies focused on reinforcement learning [34, 48, 56]. This allowed robots to learn from their environments, although high sample complexity and difficult reward engineering were involved. More recent studies [10, 19, 11, 24, 31, 60] have geared towards learning policies by imitating actions in expert demonstrations, which has yielded promising results. In addition, the robotic manipulation community has been further propelled by the introduction of new benchmarks. Transitioning from benchmarks with limited scales and diversity [74, 53, 22] to larger scale benchmarks [32, 47, 63, 49] with the increased object, scene, and task variations (e.g., RLBench [32] and CALVIN [47]).
@@
Robotic visual pre-training. To efficiently teach robots manipulation skills in real-life scenarios, it is crucial to pre-train models on large-scale data. This ensures that the generalizable visual representation across different environments is effectively learned. Drawing inspiration from the representation learning [28, 26, 64] in computer vision, there have been several endeavors [5, 6, 63, 77, 40, 72] to accomplish this goal in robotic manipulation. The RT-series [5, 6, 63], for instance, developed vision-language-action models using large-scale data that encompassed out-of-domain data from the internet and self-gathered robotic demonstrations. RoboFlamingo [40] learned to generate robot actions by fine-tuning pre-existing vision-language foundation models on robot demonstrations. However, the collection of large-scale robot demonstration data proved costly, and unresolved issues remain regarding the different data distributions between different embodiments [9, 72, 63].
@@
In parallel, several studies [50, 68, 13, 24, 52, 33, 44, 36, 45, 46] have proposed methods to learn a generalizable representation from large-scale human datasets, such as Ego4D [21] and Kinetics [8]. These methods are based on the assumption that the dynamics of human-object interactions in these datasets are analogous to those of robots performing tasks. For instance, the R3M [50] model was pre-trained on the Ego4D dataset using time-contrastive learning and video-language alignment techniques. Similarly, the MVP [68] model employed masked modeling to learn visual dynamics, and the data4robotics [13] performed a dataset-centric analysis of robotic pre-training, contributing valuable insights to the field. Despite these advances, significant challenges remain due to the inherent morphological differences between humans and robots. These differences lead to a human-robot domain discrepancy [29, 41, 58, 30, 1, 70], which limits the generalizability of models trained on human data when applied to downstream robotic tasks.
@@
Human-robot domain transfer. Many works [45, 55, 69, 75, 4, 1, 39] have been dedicated to learning robotic manipulation skills by imitating human actions in human videos. For large-scale pre-training using human data, some works [29, 58, 41] have noticed the human-robot domain gap problem and have developed two main strategies to address it. The first is to design more effective proxy tasks to learn semantics that can be better transferred to downstream robotic manipulation when pre-training models on the human data. For example, Huo et al. [29] proposed human-oriented multi-task fine-tuning on top of pre-trained models to encode meaningful semantics for manipulation. However, the manipulation-oriented proxy tasks are difficult to define, and this implicit fine-tuning without accessing robot data is not effective in mitigating the human-robot domain discrepancy. The second strategy [58, 41] directly fine-tunes the pre-trained models on downstream manipulation data to bridge the semantic gap between robot data and pre-trained human data. For example, Sharma et al. [58] proposed to insert lightweight adapters into pre-trained models, aiming to reduce the performance gap between frozen pre-trained representation and downstream full-parameter end-to-end fine-tuning. However, this strategy requires custom modifications to the pre-trained models for each specific environment, which will compromise the models’ generalizability.
@@ 3 Human-Robot Semantic Alignment Method
Considering the large domain discrepancy between the pre-trained human data and downstream robotic manipulation data in the existing pre-training paradigm, this work proposes a new adaptation paradigm. This paradigm leverages the semantic-aligned human-robot demonstration data to guide the adaptation of existing pre-trained models, effectively enhancing their generalization capabilities across various downstream tasks. Driven by this idea, we propose a simple yet effective Human-Robot Semantic Alignment method.
@@ 3.1 Problem formulation
Assuming we have a dataset $\mathcal{D}$ encompassing $N$ human-robot video pairs, i.e., $\mathcal{D}=\{H_{i},R_{i},L_{i}\}_{i=1}^{N}$, where $L_{i}$ is a sentence describing what task is being performed in the video pair, $H_{i}$ denotes the $i$-th human video demonstration, and $R_{i}$ denotes the $i$-th robot video demonstration that imitates the human actions in $H_{i}$ via teleoperation. Thus the dynamic semantics in each human-robot video pair are well-aligned. For learning the downstream manipulation policies, instead of using the robot’s raw image observations $I$ as state input, existing visual pre-training methods train visual encoder $\mathcal{F}$ on human data, and use the feature representation $\mathcal{F}(I)$ of robot’s observations as state input. In this work, to mitigate the human-robot domain discrepancy in the pre-trained representation of $\mathcal{F}$, we leverage the aforementioned paired human-robot video dataset $\mathcal{D}$ to adapt the pre-trained model $\mathcal{F}$. In this way, the adapted model $\mathcal{T}$ can extract more effective state representation for learning policies in downstream manipulation tasks.
@@ 3.2 Overview
Figure 2 provides an overall diagram of our Human-Robot Semantic Alignment (HR-Align) method. Taking the paired human and robot videos as input, the HR-Align method utilizes a frozen pre-trained model to obtain the feature representations of human video and robot video, respectively. However, since the used backbone model is pre-trained on human data, the extracted feature semantics of human and robot videos suffer from domain discrepancy. For the robot video, our HR-Align injects learnable adapter modules into the pre-trained model, so that the pre-trained model can be adapted to effectively encode the dynamic semantics in the robot video. Besides, for each stream, the corresponding task description of the videos will serve as the query to extract task-aware semantics from the aforementioned feature representations. To guarantee the introduced adapter modules can be effectively trained for adapting the pre-trained model to the robot domain, the HR-Align leverages a contrastive alignment loss, which constraints the semantics in robot video encoded by the adapted model to align with the pre-trained feature semantics in the paired human video.
3.3 HR-Align Learning
@@
where $\Upsilon$ denotes the frame sampling operation, and $h^{f},r^{f}\in\mathbb{R}^{\mathrm{\hat{T}\times\hat{H}\times\hat{W}\times\hat{C}}}$ are the frozen features extracted from the human video and robot video, respectively. Since the backbone model $\mathcal{F}$ is well pre-trained on large-scale human data, the visual dynamic semantics in human video $H$ can be effectively captured. However, due to the significant human-robot domain gap, it is difficult to extract corresponding semantics from the paired robot video, when using the human-data trained model $\mathcal{F}$ that does not access any knowledge in the robot domain during pre-training.
@@
Adapted robot stream. To this end, our HR-Align method aims to efficiently adapt the pre-trained model on the robot videos, such that the newly adapted pre-trained model $\mathcal{T}$ can alleviate the domain discrepancy. To achieve this, inspired by existing parameter-efficient fine-tuning approaches [73, 51], we form the adapted pre-trained model $\mathcal{T}$ by inserting learnable adapter modules (Eq. (3)) into the pre-trained model $\mathcal{F}$. Then we apply this model to the robot video as follows:
$r^{t}=\mathcal{T}(\Upsilon(R))=\mathcal{F}_{\textit{Adapter}}(\Upsilon(R)),$ (2)
@@
where $r^{t}\in\mathbb{R}^{\mathrm{\hat{T}\times\hat{H}\times\hat{W}\times\hat{C}}}$ is the adapted robot video features. The adapter modules can be inserted into any position of the pre-trained model $\mathcal{F}$. Without loss of generality, given the output $r^{f,inter}\in\mathbb{R}^{\mathrm{\bar{T}\times\bar{H}\times\bar{W}\times\bar{C}}}$ of an intermediate layer of model $\mathcal{F}$, the adapter module can be applied as follows:
$r^{t,next}$ $=\textit{Adapter}(r^{f,inter})$ (3)
$=r^{f,inter}+\text{Conv}_{up}(g(\text{Conv}_{down}(r^{f,inter}))),$
@@
where the residual output $r^{t,next}$ will be processed by subsequent layers of the model $\mathcal{F}$. Following existing practice [73, 51], our adapter module includes an activation function $g$, and two convolution layers (i.e., $\text{Conv}_{down}$ and $\text{Conv}_{up}$) used for feature down-projection and up-projection, respectively.
@@
Task-aware feature modeling. With the extracted frozen human feature $h^{f}$, frozen robot feature $r^{f}$, and the adapted robot feature $r^{t}$, we incrementally contribute a feature enhancement module to obtain task-aware feature representations of human and robot videos, where the corresponding task description $L$ serves as the query. The feature of task description query $l\in\mathbb{R}^{\hat{C}}$ is obtained using a frozen BERT model [15, 54], with a learnable fully connected layer followed to match the channel dimension of the video features. The process is formulated as $l=\textit{Linear}(\textit{Bert}(L))$. Later, an attention-based aggregation is defined, where the task description feature $l$ serves as the query, and the video features in each stream serve as keys and values. Taking the reshaped video features $r^{t}\in\mathbb{R}^{\mathrm{(\hat{T}\cdot\hat{H}\cdot\hat{W})\times\hat{C}}}$ in the adapted robot stream as an example, the task-aware adapted robot feature $\bar{r}^{t}$ is defined as:
@@
For the frozen human and robot data stream, the task-aware frozen human feature $\bar{h}^{f}\in\mathbb{R}^{\hat{C}}$ and task-aware frozen robot feature $\bar{r}^{f}\in\mathbb{R}^{\hat{C}}$ can be obtained in the same way. While not a core contribution, the ablation study in the appendix demonstrates that the feature enhancement effectively improves the model’s performance.
Human-Robot contrastive alignment. To mitigate the human-robot domain discrepancy in human-data pre-training, our key idea is to finetune the pre-trained model with a parameter-efficient design, where the semantic-aligned human-robot video pairs serve as bridges for the adaptation. For each human-robot video pair, the human-data pre-trained model can effectively encode the visual semantics in the human video. To guarantee the consistent semantics in the paired robot video can be captured by the adapted model, we propose a human-robot contrastive alignment loss to modulate the adaptation of the pre-trained model.
@@
In Eq. (6), $\mathcal{S}$ is a similarity function defined as $\mathcal{S}(x,y)=\exp(x^{T}y/\tau)$, where $\tau$ is a temperature factor. Under the constraint of the contrastive alignment loss, the learnable adapter modules enable the pre-trained model to capture robotic dynamics, thereby simultaneously mitigating the domain discrepancy and avoiding the need to customize models for each downstream robotic environment.